What is the efficiency of a heat engine that has an input tmperature of 500 degrees celsius and an exhaust of 250 degrees celsius? A. 50% B. 32 % C. 22%.
Answer #1

Solution:- Input temperature (T_1) :- 500 degrees

Output temperature (T_2) :- 250 degrees

Efficiency, \eta = (1 - T_2/T_1)*100% = (1 - 250/500)*100% = 50%.

Hence, efficiency = 50%. Option (A)